OpenStudy (anonymous):

Solve 2sin2x-sqrt (2) =0 in the interval [0, 2pi)

OpenStudy (anonymous):

\[ 2 \sin(2x) = \sqrt 2\\ \sin(x)= \frac{\sqrt 2}2 \]

OpenStudy (anonymous):

What angles in \[ [0, 2\pi) \] have sin equal to \[ \frac{ \sqrt2} 2\]?

OpenStudy (jdoe0001):

$$ 2 \sin(2x) = \sqrt 2 \implies \sin(2x)= \cfrac{\sqrt 2}{2}\\ sin^{-1}(\sin(2x)) = sin^{-1}\pmatrix{\cfrac{\sqrt 2}{2}}\\ 2x= \color{red}{\alpha} \implies x=\cfrac{\color{red}{\alpha}}{2} $$ as @eliassaab said, what are those angles?

OpenStudy (anonymous):

\[ \frac \pi 4 \\ \frac { 3 \pi} 4 \]

OpenStudy (anonymous):

Sory, I thought you were asking me?

OpenStudy (jdoe0001):

hehe, :), anyhow, he's not around for now, he'll be back :)